OpenStudy (anonymous):

–6p + 7 = 3(2p – 3) – 4(–10 + 4p) p = 6 p = 5 p = 7 p = 12

OpenStudy (anonymous):

\[-6p + 7 = 6p - 9 + 40 -16p\]

OpenStudy (anonymous):

\[p=\]\[4p=24\]\[4p + 7=31\]\[-6p + 7 = -10p + 31\]

OpenStudy (anonymous):

\[p = 6\]

OpenStudy (anonymous):

how did you get 4p = 24 and such ? thanks :D

OpenStudy (anonymous):

sorry i didnt know how to use this thing lol it is backwards so... -6p + 7 = 6p - 9 + 40 - 16p -6p + 7 = -10p + 31 (combining like terms) 4p + 7 = 31 (adding 10p to both sides) 4p=24 (subtracting 7 from both sides) p=6 (dividing by 4 to isolate the variable, on both sides)

OpenStudy (anonymous):

thank you you helped so much !

OpenStudy (anonymous):

No Problem :D and you are WELCOME!!!

OpenStudy (anonymous):

that was helpful. tanks so much